Let R be the relation on Z

 (the set of integers) defined by 

(x, y)  R iff x

2 + y2 = 2k for some integers k  0.

Answer questions 13 to 15 by using the given relation R.

Question 13

Which one of the following is an ordered pair in R?

1. (1, 0)

2. (2, 9)

3. (3, 8)

4. (5, 7)

Question 14

R is symmetric. Which one of the following is a valid proof showing that R is symmetric?

1. Let x, y  Z be given.

Suppose (x, y)  R

then x

2 + y2 = 2k for some k  0.

ie y

2 + x2 = 2k for some k  0.

thus (x, y)  R.

2. Let x, y  Z be given.

Suppose (x, y)  R

then x

2 + y2 = 2k for some k  0.

ie y

2 + x2 = 2k for some k  0.

thus (y, x)  R.

3. Let x, y  Z

 be given.

Suppose (x, y)  R

then x

2 + y2 = 2k for some k  0.

thus (y, x)  R.

4. Let x, y  Z be given.

Suppose (x, x)  R

then x

2 + y2 = 2k for some k  0.

ie y

2 + x2 = 2k for some k  0.

thus (y, y)  R.



1
Expert's answer
2021-07-12T13:13:01-0400

Let "R" be the relation on "\Z" defined by "(x, y)\in R" iff "x^2 + y^2 = 2k" for some integer "k\ge0."


Question 13


Taking into account that "1^2+0^2=1, 2^2+ 9^2=85, 3^2+ 8^2=73" are odd, and "5^2+ 7^2=74=2\cdot 37" is even, we conclude that only "(5,7)\in R."


Answer: 4




Question 14


Valid proof is the following:


2. Let "x, y\in\Z" be given. Suppose "(x, y) \in R", then "x^2 + y^2 = 2k" for some "k \ge 0", i.e. "y^2 + x^2 = 2k" for some "k\ge 0". Thus "(y, x) \in R."


Answer: 2